Section 13  Combinatorics

Section 13: Combinatorics

Invited Lectures

Saturday, 22 August  H 2032

14:00 – 14:45 Béla Bollobás, University of Memphis, Memphis, USA and University of Cambridge, Cambridge, UK Hereditary properties of graphs: asymptotic enumeration, global structure, and colouring (I: p.153) Chair: Hans J. Prömel

15:00 – 15:45 Harald Niederreiter, Austrian Academy of Sciences, Vienna, Austria Nets, (t, s)-sequences, and algebraic curves over finite fields with many rational points

(I: p.159)

Chair: Dieter Jungnickel

Monday, 24 August  H 2032

14:00 – 14:45 András Frank, Eötvös Lorand University, Budapest, Hungary Applications of Relaxed Submodularity (I: p.155) Chair: Alexander Schrijver

15:00 – 15:45 Jiří Matoušek, Charles University, Praha, Czech Republic Mathematical snapshots from the computational geometry landscape (I: p.158) Chair: Jaroslav Nešetril

Tuesday, 25 August  H 2032

16:15 – 17:00 Neil J.A. Sloane, AT&T Research Labs, Florham Park, USA Recent developments on the sphere-packing problem (I: p.161) Chair: Martin Aigner

17:15 – 18:00 Joseph Adolphe François Thas, University of Gent, Gent, Belgium Finite Geometries, Varieties and Codes (I: p.161) Chair: Theo Grundhöfer

Wednesday, 26 August  H 2032

14:00 – 14:45 Alain Lascoux, CNRS, Université de Marne La Vallée, Noisy-le-Grand, France Ordering the symmetric group: Why use Iwahori-Hecke algebras? (I: p.157) Chair: Andrei Zelevinsky

15:00 – 15:45 Andrei V. Zelevinsky, Northeastern University, Boston, USA Multisegment duality, canonical bases and total positivity (I: p.163) Chair: Mikhail M. Kapranov

Combinatorics  Section 13

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Geometry I  H 106

Chair: Jörg Rambau

16:15 – 16:30 Dietmar Cieslik, University of Greifswald, Greifswald, Germany The Steiner Ratio of Metric Spaces (p.275)

16:30 – 16:45 Michael Joswig, FB Mathematik, TU, Berlin, Germany On Elation Laguerre Planes (p.277)

16:45 – 17:00 Alexandre A. Makhnev, Institute of Mathematics and Mechanics, Ekaterinburg, Russia Partial geometries and the Krein condition (p.282)

Geometry II  H 106

Chair: Günter M. Ziegler

17:15 – 17:30 Min Yan, Hong Kong University, Hong Kong, China, Peoples Republic of Stratified Approach in Combinatorics (p.287)

17:30 – 17:45 Jörg Rambau, Konrad-Zuse-Zentrum für Informationstechnik, Berlin, Germany A new View on the Cayley Trick (p.283)

17:45 – 18:00 Masahiro Hachimori, University of Tokyo, Tokyo, Japan Non-constructible triangulations of balls (p.277)

Short Communications  Thursday, 20 August
Combinatorics of Words  H 106

Chair: Emo Welzl

16:15 – 16:30 Julien Cassaigne, Institut de Mathématiques de Luminy, Marseille, France On the conjectured subword complexity of the Kolakoski sequence (p.275)

16:30 – 16:45 Martin Klazar, Universität Bonn, Bonn, Germany Combinatorics of a containment relation that generalizes Davenport–Schinzel sequences (p.279)

16:45 – 17:00 Alexander Sergeevich Makarenko, National Technical University of Ukraine, Kiev, Ukraine The complexity measure of individual geometrical and algebraical objects (p.281)

Independent Sets and Coloring  H 106

Chair: Heinz Adolf Jung

17:15 – 17:30 Stephan Brandt, Freie Universität, Berlin, Germany Pach’s Theorem on triangle-free graphs revisited (p.274)

17:30 – 17:45 Kevin A. Broughan, University of Waikato, Hamilton, New Zealand Computing the minimal colouring of large graphs (p.274)

17:45 – 18:00 Ebadollah S. Mahmoodian, Sharif University of Technology, Tehran, Iran Defining sets in graph colorings (p.281)

Short Communications  Friday, 21 August
Paths and Cycles  H 106

Chair: Stephan Brandt

16:15 – 16:30 Angeles Carmona, Universitat Politècnica de Catalunya, Barcelona, Spain On the Shortest Paths in a Distance Regular Graph (p.275)

16:30 – 16:45 Min Aung, Assumption University, Bangkok, Thailand A lower bound for the circumference of a graph in terms of connectivity and minimum degree

(p.283)

16:45 – 17:00 Henda Swart, University of Natal, Durban, South Africa Generalised Measures of Distance in Graphs (p.286)

Section 13  Combinatorics

Short Communications  Saturday, 22 August
Geometry III  H 106

Chair: Klaus Altmann

16:15 – 16:30 Jennifer Mc Nulty, The University of Montana, Missoula, MT, USA Connected Hyperplanes in Binary Matroids (p.282)

16:30 – 16:45 Andreas W.M. Dress, Universität Bielefeld, Bielefeld, Germany Highly Regular Spatial Structures: Theory, Software, and Visualization (p.276)

16:45 – 17:00 Klaus Altmann, Humboldt–Universität, Berlin, Germany Hodge numbers for polytopes (p.274)

Codes and Cryptography  H 106

Chair: A. Donald Keedwell

17:15 – 17:30 Gerhard R. Paseman, Prado Internet Access, Inc., Hayward, CA, USA A different approach to Hadamard’s maximum determinant problem (p.283)

17:30 – 17:45 Gholamrega B. Khosronshahi, Inst. for Studies in Theo. Physics and Math., Tehran, Iran On trades and their codes (p.279)

17:45 – 18:00 A. Donald Keedwell, University of Surrey, Guildford, Great Britain Crossed-inverse quasigroups and their use in cryptography (p.278)

Short Communications  Monday, 24 August
Classical Combinatorial Structures  H 106

Chair: Stefan Felsner

16:15 – 16:30 Mario-Osvin Pavcevic, University of Zagreb, Zagreb, Croatia On two series of symmetric designs (p.283)

16:30 – 16:45 Christian Krattenthaler, Universität Wien, Wien, Austria New results in the enumeration of rhombus tilings (p.280)

16:45 – 17:00 Martin Kochol, FB Mathematik, FU, Berlin, Germany Superposition — a method for constructing graphs without nowhere-zero k-flows (p.280)

Matrices, Posets, and Lattices  H 106

Chair: Jens Gustedt

17:15 – 17:30 Konrad Engel, Universität Rostock, Rostock, Germany An upper bound for the size of a largest antichain in the poset of partitions of an integer (p.276)

17:30 – 17:45 Antoni Smoluk, Wroclaw University of Economics, Wroclaw, Poland Fix-point property of finite partially ordered sets (p.284)

17:45 – 18:00 Christian Bey, Universität Rostock, Rostock, Germany An intersection theorem for the function lattice (p.274)

Short Communications  Tuesday, 25 August
Subgraphs  H 106

Chair: Béla Bollobás

14:00 – 14:15 Oliver M Riordan, SFB 343, Universität Bielefeld, Bielefeld, Germany Spanning subgraphs of random graphs (p.284)

14:15 – 14:30 Diane M. Spresser, James Madison University, Arlington, Virginia, USA Graphs with degree sets of maximal cardinality II (p.285)

14:30 – 14:45 Andrew G. Thomason, University of Cambridge, Cambridge, Great Britain Hereditary hypergraph properties (p.286)

Combinatorics  Section 13

Finite Groups and Cayley Graphs  H 106

Chair: Andreas W.M. Dress

15:00 – 15:15 Reinaldo Giudici, Universidad Simón Bolívar, Caracas, Venezuela Cayley’s Graphs of order 12 (p.276)

15:15 – 15:30 Cai-Heng Li, Mathematics Dept. UWA, Nedlands, WA, Australia On Isomorphisms of Finite Cayley Graphs (p.280)

15:30 – 15:45 Mikhail H. Klin, Ben-Gurion-University of the Negev, Beer-Sheva, Israel On transitive permutation groups without semi-regular subgroups (p.279)

Short Communications  Wednesday, 26 August
Combinatorial Constructions  H 106

Chair: Mark Kayll

16:15 – 16:30 Abdelkader Khelladi, USTHB Institut de Mathematiques, El Alia, Alger, Algeria Some Results on the Fiber Product of Graphs (p.278)

16:30 – 16:45 Andrew J. Woldar, Villanova University, Villanova, PE, USA General properties of algebraically defined graphs (p.287)

16:45 – 17:00 Mark Kayll, The University of Montana, Missoula, MT, USA Stochastic independence properties of hard-core distributions (p.278)

Geometry IV  H 106

Chair: Yuri Grigoriyevich Stoyan

17:15 – 17:30 Alexei A. Tuzhilin, Moscow State University, Moscow, Russia Planar Local Minimal Networks with Convex Boundaries (p.286)

17:30 – 17:45 Alexandr O. Ivanov, Moscow State University, Moscow, Russia Linear Networks in Euclidean Spaces R^{n} (p.277)

17:45 – 18:00 Yuri Grigoriyevich Stoyan, Inst. for Problems in Machinery, Kharkov, Ukraine Properties of combinatorial set images into the arithmatic Euclidean spaces (p.285)

Poster Session  Friday, 21 August 14:00 – 15:45  MA 144

André Kündgen, Universität Ulm, Ulm, Germany

Turán problems for weighted graphs (p.280) Frank H. Lutz, Technische Universität Berlin, Berlin, Germany

Combinatorial manifolds with vertex-transitive automorphism group (p.281) Slobodan K. Simić, University of Belgrade, Belgrade, Yugoslavia

On graphs whose second largest eigenvalue does not exceed the golden section (p.284) Indulis Strazdins, Riga Technical University, Jelgava, Latvia

Universal classification of Boolean functions by the affine and linear groups (p.285) Kai-Yao Xiang, No. 1 High School, Dongfeng Motor Corporation, Hamamatsu, Japan

The Minimal Group with Graph and Its Permutation Elements (p.287) Tomoyuki Yoshida, Hokkaido University, Sapporo, Japan

Exponential formulas and Krull-Schmidt categories (p.288)